OpenStudy (anonymous):

integration : x(e^(3x)) dx

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

i tried to use 'u' sub but it didn't work.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

Use integration by parts. \[u = x \\ dv = e^{3x} dx\]

OpenStudy (anonymous):

it can't be used isn't it? since dv doesn't have 'x dx'

OpenStudy (anonymous):

You don't need an x, just a dx. And you have one.

OpenStudy (anonymous):

ok let me try it

OpenStudy (anonymous):

I'll do \(\int x e^x dx\) for you as an example. Then you should see how to do it. \[u = x \\ du = dx \\ dv = e^x dx \\ v = e^x\] So plugging into \[uv - \int v du\] We get \[xe^x - \int e^x dx\] Resulting in: \(xe^x - e^x + C\)